### v2

# Automated Data Augmentation for Few-Shot Time Series Forecasting: A Reinforcement Learning Approach Guided by a Model Zoo

Abstract
Time series forecasting, particularly in few-shot learning scenarios, is challenging due to the limited availability of high-quality training data. To address this, we present a pilot study on using reinforcement learning (RL) for time series data augmentation. Our method, ReAugment, tackles three critical questions: which parts of the training set should be augmented, how the augmentation should be performed, and what advantages RL brings to the process. Specifically, our approach maintains a forecasting model zoo, and by measuring prediction diversity across the models, we identify samples with higher probabilities for overfitting and use them as the anchor points for augmentation. Leveraging RL, our method adaptively transforms the overfit-prone samples into new data that not only enhances training set diversity but also directs the augmented data to target regions where the forecasting models are prone to overfitting. We validate the effectiveness of ReAugment across a wide range of base models, showing its advantages in both standard time series forecasting and few-shot learning tasks.
1 Introduction
Time series forecasting is a critical task with diverse applications, but it faces significant challenges due to the limited availability of high-quality training data. This challenge is further amplified in time-evolving domains with inherent non-stationarity and becomes even more pronounced in few-shot learning scenarios, where the scarcity of data severely limits the performance of forecasting models. While recent advancements have focused on developing more effective deep-learning architectures to capture long-term trends, clinical patterns, and multivariate relationships in the data (Wu et al., 2021; Nie et al., 2023; Liu et al., 2024b), in this work, we explore a data-centric approach and introduce a novel, learning-based data augmentation method that can be seamlessly integrated with existing forecasting models.
We recognize that effective data augmentation requires generating high-quality, diverse training samples. However, in practice, existing forecasting models typically rely on fixed-form data augmentation techniques (Wen et al., 2021; Cheung & Yeung, 2020), which lack data-dependent adaptability and may introduce unexpected noises. Previous learning-based augmentation methods typically involve contrastive learning (Demirel & Holz, 2024) or Mixup combinations (Schneider et al., 2024) to generate new data sequences. However, these methods are not task-specific, as the data generation process is not guided by the performance of forecasting models. In contrast, we argue that aligning the training objectives of augmentation models with the resulting forecasting performance is essential, enabling a closed-loop optimization of the data augmentation process.
In this work, we present ReAugment, which uses reinforcement learning (RL) to create the closed-loop data augmentation process. Our goal is to address data scarcity and enhance the generalizability of forecasting models in few-shot learning scenarios. We aim to answer three critical questions: which parts of the training set should be augmented, how the augmentation should be performed, and what advantages RL offers compared to previous approaches. Specifically, ReAugment dynamically identifies a training subset of overfit-prone data samples that would benefit most from augmentation and automatically searches for optimal augmentation policies tailored to these samples.
ReAugment initially leverages a forecasting model zoo, consisting of diverse instances of the same network architecture trained via cross-validation, to identify overfit-prone training samples in need of augmentation. An interesting finding is that training forecasting models exclusively with data samples that present low prediction variance across the model zoo often leads to better performance (detailed in Section 3). This insight leads us to select the data points with high prediction variance across the model zoo as the root of the subsequent augmentation process. Additionally, targeting overfit-prone samples for augmentation ensures more efficient use of limited data resources, compared to applying uniform augmentation to all training samples.
Based on the identified overfit-prone samples, ReAugment learns an augmentation network in forms of a variational masked autoencoder (VMAE), where we use the latent space as the action space to be optimized in the RL framework. The RL-based augmentation policies are guided by a reward function derived from the backtesting errors, measured by applying the generated data to the forecasting model zoo. The training objective balances both the diversity and quality of the augmented data: It encourages the new data points to fall into regions where the forecasting models tend to overfit, while maintaining alignment with the original data distribution. Specifically, we employ the REINFORCE algorithm (Williams, 1992) to enable the backtesting diversity evaluated on the model zoo, which is non-differentiable, to guide the learning process of the augmentation network.
The contributions of this work are as follows:
Key idea: We propose ReAugment, which offers a pilot study on leveraging RL for time series augmentation.
Finding anchor data: We propose a novel method for identifying overfit-prone samples by leveraging cross-validation errors from a forecasting model zoo to pinpoint anchor points with potential value for data augmentation.
Model: We introduce a novel network architecture based on a variational masked autoencoder for sequential data, where the latent space is used as the action space of RL.
Algorithm: We define a specific reward function that incorporates the model zoo’s prediction responses to guide the generation of new data points targeting regions where the forecasting models are prone to overfitting.
Results: ReAugment demonstrates substantial improvements across a wide range of base models and datasets, especially in few-shot learning scenarios.
2 Related Work
Transformer-based time series forecasting.
Compared with CNN-based or RNN-based forecasting models (Torres et al., 2021; Wang et al., 2022; Che et al., 2018; Sagheer & Kotb, 2019), recent Transformer-based methods (Li et al., 2019; Wu et al., 2021; Zhou et al., 2021; Liu et al., 2021; Zhou et al., 2022; Zhang & Yan, 2022; Zeng et al., 2023; Cao et al., 2024; Yi et al., 2024) have shown superior performance across a wide range of time series forecasting tasks. Most previous work has focused on developing more effective network architectures to capture long-term trends, clinical patterns, and multivariate relationships within the data. For instance, Autoformer (Wu et al., 2021) improves upon the Transformer by employing a deep decomposition architecture that progressively separates trend and seasonal components throughout the forecasting process. PatchTST (Nie et al., 2023) vectorizes time series data into patches of specified size, which are then encoded through a Transformer, with the model producing forecasts of the desired length via an appropriate prediction head. LSTF-Linear (Zeng et al., 2023) simplifies complex time series forecasting problems and outperforms many Transformers by using a set of remarkably simple one-layer linear models. iTransformer (Liu et al., 2024b) modifies the architecture by adopting components with inverted dimensions, demonstrating superior performance on multivariate time series data. Unlike these methods, we explore a data-centric approach to time series forecasting, rather than focusing on the model architecture. We introduce a data augmentation method that can be seamlessly integrated with existing forecasting models.
Few-shot time series forecasting.
Few-shot time series forecasting refers to the ability to make accurate predictions for time series data with very limited historical information. Traditional time series forecasting models, such as ARIMA, Exponential Smoothing, and deep learning models like LSTMs or the above Transformer-based models, generally require large amounts of historical data to make reliable predictions. However, in many real-world applications, obtaining sufficient data can be challenging, especially in scenarios where data is scarce, noisy, or difficult to collect (Dooley et al., 2023; Xu et al., 2024; Jiang et al., 2023; Yuan et al., 2024). Notably, recent literature has introduced large foundation models specifically designed for time series forecasting (Das et al., 2024; Jin et al., 2024; Bian et al., 2024; Ekambaram et al., 2024; Liu et al., 2024a, c; Pan et al., 2024), often evaluating these models under zero-shot domain generalization settings. However, in our preliminary experiments, we found that as the distribution gap between training and testing data increases (e.g., when data comes from entirely different domains), the generalization performance of these models significantly decreases.
Time series augmentation.
Iglesias et al. (2023) have presented a taxonomy of augmentation techniques. Simple augmentation methods involve time, frequency, and magnitude domain transformation techniques such as slicing (Cao et al., 2020), frequency warping (Cui et al., 2015), and jittering (Flores et al., 2021). The second category is the learning-based methods, including those based on contrastive learning (Demirel & Holz, 2024), and Mixup combinations (Schneider et al., 2024). Additionally, advanced generative models have been employed to generate realistic time series data, including the GAN-based (Yoon et al., 2019; Liao et al., 2020), VAE-based (Sohn et al., 2015; Li et al., 2020), and Diffusion-based (Huang et al., 2023) methods. The generated samples can be used for further training of the forecasting models. In contrast to existing approaches, we present a pilot study on using reinforcement learning for time series data augmentation.
3 Overfit-Prone Samples as Anchor Data
The first challenge we need to address is identifying which parts of the training set would benefit the most, so they can be used as anchor points in the augmentation process.
3.1 Finding Overfit-Prone Data with a Model Zoo
A key assumption in time series forecasting, particularly in few-shot learning scenarios, is that forecasting models tend to overfit certain regions of the training data. Different training samples can have varying impacts on training quality. Intuitively, we aim to generate new data with distributions around the original data points that have higher probabilities of overfitting. To achieve this, we measure the cross-validation errors from a batch of forecasting models to pinpoint the overfit-prone samples.
To construct the forecasting model zoo, we divide the training set into parts and perform -fold cross-validation. In this way, we obtain sets of model parameters for the same network architecture, e.g., iTransformer (Liu et al., 2024b), which we denote as . For each data point , we evaluate the prediction errors of the other models that were not trained on this subset, along with the training error of the model trained on this subset. We then calculate the variance of the MSE across the models, denoted as , and sort the “model-zoo variance” of all data points.
The model-zoo variance measures the inconsistency in predictions made by different models trained on slightly different subsets, and can therefore serve as an indicator of the overfit-prone samples. When a model overfits, its predictions on certain data points are highly sensitive to the specific subset it was trained on, leading to a high variability of the performance across the models. As a result, data points with high model-zoo variance are more likely to be in regions where the models struggle to generalize, making them overfit-prone and thus ideal candidates for data augmentation to improve generalization.
In practice, we split the training set into two subsets based on the top and bottom values of the model-zoo variance, and refer to the top subset as the overfit-prone samples.
3.2 Preliminary Findings on Overfit-Prone Data
To understand the impact of the overfit-prone samples on the training process, we conduct the following experiments. We train forecasting models separately using data from Group A (the subset with top model-zoo variance) and Group B (the subset with bottom model-zoo variance). We evaluate the models on the same test set. Given its superior average performance, we select iTransformer as the preferred model for these experiments. The experiments are conducted on classic real-world multivariate time series benchmarks, including ETT, Weather, and Electricity.
The forecasting results are shown in Figure 2, where a lower MAE indicates more accurate predictions. Our experiments demonstrate that the choice of training sets has a significant impact on final performance. Predictions trained in Group B consistently outperform those trained in Group A across all benchmarks, with substantial margins. This finding suggests that the overfit-prone samples are more likely to negatively affect the training quality of forecasting models. Since data augmentation is an effective strategy to mitigate overfitting, we propose generating new data around the distributions of the overfit-prone samples, allowing the forecasting model to learn more generalizable patterns from these data points.
4 ReAugment
4.1 Overall Training Pipeline
Based on preliminary findings regarding overfit-prone data, we recognize the importance of identifying such samples and using them as anchor points for data augmentation. This approach can help prevent the model from overfitting to specific patterns. By doing so, our method effectively tackles the challenges of data scarcity, leading to improved generalization performance in few-shot learning scenarios. The training pipeline of our approach consists of three stages:
Stage A: Train a probabilistic generative model to initialize the neural augmentor using overfit-prone samples. Implemented as a VMAE, the model takes partially masked time series data and corresponding absolute timestamps in the dataset as inputs to reconstruct the complete data.
Stage B: Finetune the VMAE using an RL algorithm, enabling it to generate augmented data that goes beyond merely replicating the original data distribution.
Stage C: Train the forecasting model using both the original and the augmented data.
Specifically, we augment the top overfit-prone samples, tripling the size of the original training set. This pipeline ensures that data augmentation targets overfit-prone samples, addressing their weaknesses and improving forecasting model performance in data-scarce scenarios.
4.2 Variational Masked Autoencoder as the RL Actor
By using a probabilistic generative model as the neural augmentor, denoted as , we can generate an infinite amount of data by learning a transformation function based on the overfit-prone data , where is the data index and is the length of the data sequence. The key is to learn an appropriate distribution of the latent variable , balancing the diversity of the augmented data with its similarity to the original data. For simplicity, we omit the data index in the following notations. The initial learning step involves optimizing with a data reconstruction objective, minimizing the divergence between the masked data and the original data distribution.
We design a Variational Masked Autoencoder (VMAE), as illustrated in Figure 3(left), which takes masked time series data with absolute timestamp in the whole dataset as input and outputs the complete corresponding data. The entire architecture contains four modules, including (i) the prior network, which learns the prior distribution of based on the masked data, (ii) the posterior module, which learns the posterior distribution of based on the original data, (iii) the data encoder, which extracts significant features from with , (iv) the decoder, which generates data from the encoding features and the latent variables. These modules are parametrized by and . We have
| Prior: | (1) | |||
| Posterior: | ||||
| Encoder: | ||||
| Decoder: |
We draw the latent variables, which control the diversity of the generated data, from parametrized Gaussian distributions. The mean and standard deviation of these distributions are produced by the prior and posterior modules. We minimize the Kullback–Leibler (KL) divergence between the prior and posterior distributions. After this training stage, we replace the posterior with the prior as input to the decoder. The overall objective function is
where is the set of the overfit-prone data obtained by measuring the prediction diversity over the model zoo. In line with previous work, we use the loss to measure the reconstruction error. Notably, the posterior module is used exclusively to constrain the prior learner and is not utilized in subsequent training stages.
A key aspect of our method is how we handle the absolute timestamps during the training and data augmentation phases, respectively. While VMAE is initially trained using the original absolute timestamps from the data, during the data augmentation phase in Stages B&C, we modify these timestamps by sampling them from the test set’s time range, rather than from the training time range. Empirically, this technique allows us to generate augmented samples that are more closely aligned with the distribution of the test set.
In general, the VMAE design can be integrated into any encoder-decoder-based time series forecasting architecture. In this work, we specifically adopt the encoder and decoder (implemented as a linear projector) from iTransformer (Liu et al., 2024b) for feature extraction and decoding.
4.3 REINFORCE Guided by Model Zoo Predictions
The fundamental idea behind training Stage B is that simply generating data by approximating the distribution of the original data is insufficient. Therefore, we introduce an RL framework to expand the data distribution of overfit-prone samples, with the hope that it will naturally cover important patterns in the test set and prevent the forecasting from learning trivial solutions to fit the original few-shot data.
Intuitively, we want the augmented data to have two properties: (i) it should be reasonably diversified to increase the data size and broaden the distribution around overfit-prone samples, and (ii) it should not deviate too far from the original distribution, as excessive divergence may introduce undesired noise. To achieve the first goal, inspired by our preliminary findings, a straightforward approach is to augment the dataset with samples that exhibit diverse prediction results across the model zoo. The insight is to augment the overfit-prone data with the potential overfit-prone data. The model zoo, denoted by , consists of instances of the same network architecture, obtained through cross-validation with pretrained parameters .
However, a practical challenge is that the “model-zoo variance” is not differentiable and thus cannot be directly used to optimize the augmentor through gradient descent. To tackle this problem, we propose to finetune the neural augmentor using a REINFORCE algorithm, taking the latent space generated by the prior module in VMAE as the action space. As illustrated in Figure 3, we incorporate the model-zoo variance evaluated on the augmented data as part of the reward function:
where represents the prediction results from the -th pretrained instance. Subsequently, we finetune the VMAE’s prior module through REINFORCE, treating the prior learner as the policy network while keeping the other network parameters fixed. The learning objective is to maximize the following reward function:
The reward function encourages augmented samples to exhibit more pronounced predictive variance across the model zoo, while also constraining the differences between the augmented and original data. In the reward function, a scaled-sigmoid function is employed to minimize the likelihood of rewards clustering around or controlled by hyperparameter . This approach ensures that, despite potential order-of-magnitude differences in reconstruction error and backtest variance within the model zoo, the reward function can learn effectively from these variations.
Based on this reward function, the policy network (i.e., the prior network in VMAE) is optimized as follows via gradient ascent, where is the learning rate:
5 Experiments
| Training Data | iTransformer (2024b) | PatchTST (2023) | DLinear (2023) | Average | ||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 (Raw) | 0.434 | 0.411 | 0.458 | 0.446 | 0.435 | 0.408 | 0.442 | 0.422 |
| ReAugment | 0.422 | 0.403 | 0.440 | 0.429 | 0.422 | 0.388 | 0.428 | 0.407 |
| Promotion | 2.76% | 1.95% | 3.93% | 3.81% | 2.99% | 4.90% | 3.23% | 3.55% |
| ETTh2 (Raw) | 0.362 | 0.320 | 0.367 | 0.321 | 0.402 | 0.356 | 0.377 | 0.332 |
| ReAugment | 0.339 | 0.302 | 0.349 | 0.306 | 0.369 | 0.334 | 0.352 | 0.310 |
| Promotion | 6.35% | 5.63% | 4.90% | 4.67% | 8.21% | 6.18% | 6.49% | 5.49% |
| ETTm1 (Raw) | 0.440 | 0.470 | 0.428 | 0.457 | 0.442 | 0.471 | 0.437 | 0.466 |
| ReAugment | 0.410 | 0.436 | 0.403 | 0.433 | 0.431 | 0.462 | 0.415 | 0.444 |
| Promotion | 6.82% | 7.23% | 5.84% | 5.25% | 2.49% | 1.91% | 5.05% | 4.80% |
| ETTm2 (Raw) | 0.282 | 0.204 | 0.276 | 0.199 | 0.303 | 0.219 | 0.287 | 0.207 |
| ReAugment | 0.275 | 0.196 | 0.268 | 0.193 | 0.297 | 0.216 | 0.280 | 0.202 |
| Promotion | 2.48% | 3.92% | 2.90% | 3.02% | 1.98% | 1.37% | 2.45% | 2.77% |
| Weather (Raw) | 0.231 | 0.187 | 0.232 | 0.189 | 0.277 | 0.212 | 0.247 | 0.196 |
| ReAugment | 0.229 | 0.185 | 0.227 | 0.186 | 0.276 | 0.212 | 0.244 | 0.194 |
| Promotion | 0.87% | 1.07% | 2.16% | 1.59% | 0.36% | 0.00% | 1.13% | 0.89% |
| Electricity (Raw) | 0.258 | 0.168 | 0.295 | 0.200 | 0.307 | 0.215 | 0.287 | 0.194 |
| ReAugment | 0.254 | 0.165 | 0.275 | 0.187 | 0.305 | 0.216 | 0.278 | 0.189 |
| Promotion | 1.55% | 1.79% | 6.78% | 6.50% | 0.65% | -0.47% | 2.99% | 2.61% |
| Traffic (Raw) | 0.318 | 0.466 | 0.327 | 0.541 | 0.452 | 0.724 | 0.366 | 0.578 |
| ReAugment | 0.293 | 0.429 | 0.314 | 0.521 | 0.406 | 0.663 | 0.338 | 0.538 |
| Promotion | 7.86% | 7.94% | 3.98% | 3.70% | 10.18% | 8.43% | 7.34% | 6.69% |
| Exchange (Raw) | 0.228 | 0.103 | 0.226 | 0.103 | 0.226 | 0.104 | 0.227 | 0.103 |
| ReAugment | 0.224 | 0.097 | 0.223 | 0.098 | 0.224 | 0.099 | 0.224 | 0.098 |
| Promotion | 1.75% | 5.83% | 1.33% | 4.85% | 0.88% | 4.81% | 1.32% | 5.16% |
| Dataset | Original | Gaussian | Convolve | TimeGAN | ADA | ReAugment | Promotion | |||||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.434 | 0.411 | 0.437 | 0.416 | 0.441 | 0.417 | 0.444 | 0.419 | 0.435 | 0.413 | 0.4220.01 | 0.4030.01 | 2.76% | 1.95% |
| ETTh2 | 0.362 | 0.320 | 0.365 | 0.321 | 0.364 | 0.323 | 0.366 | 0.327 | 0.368 | 0.331 | 0.3390.01 | 0.3020.01 | 6.35% | 5.63% |
| ETTm1 | 0.440 | 0.470 | 0.438 | 0.469 | 0.426 | 0.479 | 0.430 | 0.483 | 0.429 | 0.484 | 0.4100.01 | 0.4360.01 | 6.82% | 7.23% |
| ETTm2 | 0.282 | 0.204 | 0.283 | 0.204 | 0.286 | 0.207 | 0.285 | 0.206 | 0.284 | 0.207 | 0.2750.02 | 0.1960.01 | 2.48% | 3.92% |
| Weather | 0.231 | 0.187 | 0.240 | 0.196 | 0.253 | 0.204 | 0.239 | 0.191 | 0.246 | 0.198 | 0.2290.00 | 0.1850.00 | 0.87% | 1.07% |
| Electricity | 0.258 | 0.168 | 0.263 | 0.170 | 0.262 | 0.170 | 0.267 | 0.177 | 0.265 | 0.171 | 0.2540.01 | 0.1650.01 | 1.55% | 1.79% |
| Traffic | 0.318 | 0.466 | 0.319 | 0.467 | 0.320 | 0.463 | 0.315 | 0.449 | 0.318 | 0.456 | 0.2930.01 | 0.4290.01 | 7.86% | 7.94% |
| Exchange | 0.228 | 0.103 | 0.229 | 0.104 | 0.226 | 0.099 | 0.226 | 0.100 | 0.235 | 0.116 | 0.2240.00 | 0.0970.00 | 1.75% | 5.83% |
| Dataset | Original | Guassian | Convolve | TimeGAN | ReAugment | |||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.405 | 0.387 | 0.4070.02 | 0.3920.01 | 0.416 | 0.399 | 0.409 | 0.390 | 0.3960.01 | 0.3810.01 |
| ETTh2 | 0.350 | 0.301 | 0.3520.01 | 0.3070.01 | 0.356 | 0.303 | 0.348 | 0.299 | 0.3460.01 | 0.2940.01 |
| ETTm1 | 0.377 | 0.341 | 0.3740.02 | 0.3400.02 | 0.387 | 0.352 | 0.392 | 0.357 | 0.3640.01 | 0.3280.01 |
| ETTm2 | 0.272 | 0.186 | 0.2720.01 | 0.1870.00 | 0.275 | 0.188 | 0.279 | 0.190 | 0.2630.01 | 0.1790.00 |
| Weather | 0.219 | 0.178 | 0.2270.01 | 0.1870.01 | 0.265 | 0.210 | 0.219 | 0.177 | 0.2060.00 | 0.1700.00 |
| Electricity | 0.239 | 0.148 | 0.2430.01 | 0.1500.00 | 0.264 | 0.170 | 0.276 | 0.183 | 0.2360.01 | 0.1470.01 |
| Traffic | 0.269 | 0.392 | 0.2690.01 | 0.3940.02 | 0.283 | 0.407 | 0.296 | 0.412 | 0.2640.01 | 0.3880.01 |
| Exchange | 0.206 | 0.086 | 0.2080.00 | 0.0870.00 | 0.210 | 0.087 | 0.210 | 0.087 | 0.2040.00 | 0.0850.00 |
5.1 Experimental Setups
Following previous work (Wu et al., 2021; Liu et al., 2024b; Nie et al., 2023; Zeng et al., 2023), we thoroughly evaluate the proposed ReAugment on five publicly available real-world datasets: ETT (including subsets), Traffic, Electricity, Weather, and Exchange. Please refer to Appendix B for more details on these datasets.
We evaluate ReAugment in two setups: the few-shot learning setup and the standard setup, which provides full access to the original training set of the above datasets.
First, in the few-shot setup, we simulate scenarios with limited training data to assess the model’s ability to handle data scarcity. Specifically, we reduce the training set size to either or of the full dataset, depending on the dataset characteristics. The few-shot training data corresponds to the earliest portion of the time series, ensuring a significant distribution shift from the test set. The validation and test sets remain consistent with those used in prior studies.
Second, in the standard setup, we follow the configuration from previous work, allowing access to the full training set while keeping the validation and test sets unchanged.
We primarily use iTransformer (Liu et al., 2024b) for the forecasting model, as it has demonstrated strong performance in standard time series forecasting tasks. We also conduct experiments with another Transformer-based method, PatchTST (Nie et al., 2023), and the linear model DLinear (Zeng et al., 2023). Unless otherwise specified, we use a model zoo consisting of cross-validation models.
We employ a fixed lookback length of time steps across all datasets and report the multivariate sequence prediction results with prediction lengths of time steps. We provide additional implementation details in Appendix C, including information on hyperparameters, hardware requirements, and computational costs.
5.2 Baseline Augmentation Methods
Gaussian augmentation.
Inspired by prior literature (Iglesias et al., 2023), we use traditional data augmentation methods that apply simple transformations, such as adding Gaussian noise to the raw data. This approach enhances the diversity of the training data by adjusting the controllable variances and means of the added Gaussian noise.
Convolve augmentation.
We employ another traditional data augmentation method based on the function in the Tsaug library (Wen & Keyes, 2019).
TimeGAN (Yoon et al., 2019).
We also compare with the TimeGAN data augmentor, which combines supervised and adversarial objective optimization. Specifically, through a learned embedding space, the network is guided to adhere to the dynamics of the training data during sampling.
ADA (Schneider et al., 2024).
We employ the Anchor Data Augmentation (ADA) method, which improves the domain-agnostic Mixup techniques. ADA uses multiple replicas of modified samples generated by Anchor Regression (AR) to create additional training examples.
| RL | ETTh1 | ETTh2 | ETTm1 | ETTm2 | Weather | Electricity | Traffic | |||||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| 0.424 | 0.404 | 0.341 | 0.304 | 0.416 | 0.441 | 0.276 | 0.196 | 0.230 | 0.186 | 0.255 | 0.166 | 0.292 | 0.430 | |
| 0.422 | 0.403 | 0.339 | 0.302 | 0.410 | 0.436 | 0.275 | 0.196 | 0.229 | 0.185 | 0.254 | 0.165 | 0.293 | 0.429 | |
| Augmented Data | ETTh1 | ETTh2 | ETTm1 | ETTm2 | Weather | |||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| 30% Top-Variance Data | 0.426 | 0.408 | 0.339 | 0.303 | 0.415 | 0.441 | 0.265 | 0.184 | 0.228 | 0.183 |
| 50% Top-Variance Data | 0.422 | 0.403 | 0.339 | 0.302 | 0.410 | 0.436 | 0.275 | 0.196 | 0.229 | 0.185 |
| All Training Data | 0.436 | 0.412 | 0.351 | 0.311 | 0.433 | 0.460 | 0.283 | 0.203 | 0.233 | 0.188 |
5.3 Results of Few-Shot Time Series Forecasting
Under the few-shot learning setup, we evaluate the effectiveness of ReAugment on various forecasting models, including iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), and DLinear (Zeng et al., 2023). As illustrated in Table 1, on most public datasets, applying our method to construct a model zoo and perform automated data augmentation consistently improves the prediction accuracy of these two popular time series forecasting approaches. This demonstrates that our method is adaptable to different forecasting models and can effectively enhance data augmentation.
In Table 2, we compare the performance of different augmentation methods. All augmentation methods use iTransformer as the forecasting model and expand the original training set by three times. Notably, ReAugment consistently outperforms all methods across all datasets, including the state-of-the-art ADA method. In most cases, however, data-agnostic augmentation methods, such as Gaussian and Convolve, result in a negative impact on performance.
5.4 Results with Full Access to Training Set
ReAugment can also be applied to standard time series forecasting scenarios, where we have full access to the entire training sets. As shown in Table 3, ReAugment delivers significant performance improvements across multiple datasets in the standard setup, highlighting its strong generalizability beyond the few-shot learning context. Other experimental details, such as the lookback and prediction lengths, are consistent with those in the few-shot setup.
5.5 Model Analyses
A new evaluation metric for augmentation.
Evaluating the improved forecasting accuracy achieved through data augmentation provides a direct measure of our method’s effectiveness. However, it does not account for the fact that the impact of data scarcity can vary significantly across different datasets. We propose a new metric that calculates the ratio of the performance promotion achieved through data augmentation in the few-shot setup, compared to the improvement obtained using the full training set. For example, when using MSE, it can be formulated as:
This metric considers the nature of the dataset and provides a more comprehensive evaluation of the value of augmentation. As shown in Table 6, a larger value indicates a greater performance improvement due to data augmentation.
| Metric | ETTh1 | ETTh2 | ETTm1 | ETTm2 | Average |
| 41.4% | 191.7% | 47.6% | 70.0% | 87.7% | |
| 33.3% | 94.7% | 26.4% | 44.4% | 49.7% | |
| Weather | Elec | Traffic | Exc | ||
| 16.7% | 21.1% | 51.0% | 18.2% | 26.8% | |
| 22.2% | 15.0% | 50.0% | 35.3% | 30.6% |
Ablation study on REINFORCE.
The proposed RL framework enables our model to optimize the distribution of the latent variable based on the backtest results across the model zoo, thereby generating augmented data that balances data diversity and similarity to the original data. However, without RL finetuning, the original VMAE model can also be used as an independent data augmentor. Accordingly, we train the iTransformer model using augmented data by the VMAE model pretrained in Stage A. As shown in Table 4, the use of RL consistently improves the prediction results in most cases, highlighting the significance of RL finetuning for improving the quality of the augmented data.
Shall we augment all training samples?
Inspired by preliminary findings, we augment the top overfit-prone samples. What would be the impact of augmenting more or fewer samples? To investigate this, we compare augmenting the top and top overfit-prone samples with augmenting the entire few-shot dataset. For consistency, the total amount of augmented data was kept three times the size of the original training set and applied to the iTransformer model. As shown in Table 5, directly augmenting the entire few-shot dataset was less effective than data-dependent augmentation, which aligns with our preliminary findings. Furthermore, augmenting different percentages of high-variance samples resulted in varying degrees of improvement, demonstrating that our data augmentation model can adaptively enhance overfit-prone samples, leading to better performance of the forecasting model.
6 Conclusions and Limitations
In this paper, we proposed ReAugment, a novel data augmentation method driven by reinforcement learning. There are two key technical contributions in methodology: First, our method automatically identifies the critical training data for augmentation, termed the overfit-prone samples, based on prediction diversity from a set of pretrained models. Second, ReAugment exploits a variational masked autoencoder in conjunction with the REINFORCE algorithm to generate new data from these overfit-prone samples. ReAugment significantly boosts forecasting performance while maintaining minimal computational overhead by leveraging a learnable policy to transform the overfit-prone samples.
One unresolved issue in this study is the reliance on multiple pretrained models, which can increase the computational complexity of the approach. While we briefly discuss the computational cost in the appendix, this limitation is important to consider for practical applications.
Impact Statement
The potential social impact of our data augmentation method for few-shot time series forecasting could be substantial, especially in areas like economics, healthcare, climate prediction, and other fields that rely on time-sensitive forecasting. Our method could dramatically improve the accuracy of forecasts in situations where data is scarce. In industries like finance or energy, where historical data might be limited or costly to gather, being able to produce more reliable predictions with minimal data would have a significant impact. This could reduce risk and uncertainty, leading to more informed decisions.
While our method could have many benefits, it also raises some ethical questions. First, if the augmentation process isn’t well-calibrated, it could introduce biases that lead to inaccurate or unfair predictions. For instance, if certain data patterns are overly amplified, predictions in specific regions or for certain demographics could be skewed. Second, in sensitive domains like healthcare, time series data augmentation could be used to generate predictions without compromising patient privacy. However, there must be a balance between improving predictive accuracy and safeguarding individual privacy.
Appendix
Appendix A Overall Training Pipeline
We present the pseudocode of the overall training pipeline in Algorithm 1.
Appendix B Dataset Details
Here is a detailed description of the five experiment datasets:
1.
ETT consists of two hourly-level datasets (ETTh) and two -minute-level datasets (ETTm). Each of them contains factors of electricity transformers including load and oil temperature from July 2016 to July 2018.
2.
Traffic is a collection of road occupancy rates measured by sensors on San Francisco Bay area freeways from January 2015 to December 2016.
3.
ECL collects hourly electricity consumption of clients from 2012 to 2014.
4.
The Weather dataset includes meteorological indicators, such as air temperature and humidity, recorded minutes from the weather station of the Max Planck Biogeochemistry Institute in 2020.
5.
The Exchange dataset records the daily exchange rates of different countries ranging from 1990 to 2016.
For the standard setup, we follow the data processing method of iTransformer (Liu et al., 2024b), dividing the dataset into training, validation, and test sets, with this partitioning strictly aligned in chronological order.
In addition, to simulate a scenario with limited training data, we propose the few-shot setup. Specifically, we reduce the training set size to either or of the full dataset, while keeping the validation and test sets the same as in the standard setup. Notably, the training data used consists of the most distant portion of the time series relative to the test set, to simulate a more challenging time series forecasting task. In Table 7, we provide the number of variables (i.e., the feature dimension at a single time point) in each dataset, the total number of time points, and the number of time points within each set of the train-validation-test partitions for both standard and few-shot setup.
| ETTh1 / ETTh2 | ETTm1 / ETTm2 | Traffic | |
| Features | 7 | 7 | 862 |
| Time points (Standard) | 14307 | 57507 | 17451 |
| Time points (Few-shot) | 8443 | 28793 | 8756 |
| Partition (Standard) | (8545, 2881, 2881) | (34465, 11521, 11521) | (12185, 1757, 3509) |
| Partition (Few-shot) | (2681, 2881, 2881) | (5751, 11521, 11521) | (3490, 1757, 3509) |
| Electricity | Weather | Exchange | |
| Features | 321 | 21 | 8 |
| Time points (Standard) | 26211 | 52603 | 7207 |
| Time points (Few-shot) | 13136 | 21071 | 3528 |
| Partition (Standard) | (18317, 2633, 5261) | (36792, 5271, 10540) | (5120, 665, 1422) |
| Partition (Few-shot) | (5242, 2633, 5261) | (5260, 5271, 10540) | (1441, 665, 1422) |
Appendix C Implementation Details
C.1 Hyperparameters
In Table 8, we provide the hyperparameter details of VMAE and REINFORCE. For the encoder and decoder, we adopt the identical hyperparameters as those employed in iTransformer.
| Notation | Hyperparameter | Description |
| 0.001 | Learning rate of REINFORCE | |
| 0.1 | Weight of KL-divergence in the VMAE loss function | |
| 96 | Time series periods length | |
| 0.01 | Parameters of scaled sigmoid | |
| 32 | Batch size for VMAE training |
C.2 Computing Resources and Computational Costs
We perform all experiments on an NVIDIA RTX 3090 GPU. The proposed data augmentation method involves training VMAE and employing the REINFORCE algorithm, and it requires backtesting across multiple models within the model zoo. These processes introduce additional computational overhead compared to basic prediction models. As shown in Table 9, we report the training time for VMAE (Stage A), REINFORCE (Stage B), and the forecasting model (Stage C) under the few-shot setup. The increased training time (including Stage A and Stage B) is considered acceptable due to the significant performance gains. Besides, it is notably shorter than the time needed for training the forecasting models.
| Dataset | Stage A: VMAE | Stage B: REINFORCE | Stage C: Forecasting | |
| iTransformer | PatchTST | |||
| ETTh1 | 1min | 2min | 2min | 5min |
| Elec. | 24min | 31min | 1h 33min | 2h 24min |
| Traffic | 1h 22min | 1h 53min | 4h 27min | 6h 50min |